A point charge q1 = +3.90 is at origin. How far should the second point charge of +6.30  be placed to have electric potential energy of 0.700 J?



1
Expert's answer
2022-03-29T12:29:29-0400

Answer

The distance must

"R=\frac{Kqq'}{U}\\=\frac{9*10^9*3.9*6.30}{0.7}\\=315.9m"